Please indicate all steps taken to solve this question clearly so it is easy to understand. Thank you.

Problem #3 : If, in 11 years, $1000 increased to $2850, what is the effective monthly rate, as a percentage? Percentage, correct to 2 decimals. (Do not include the % sign in your answer.) Problem #3: 0.79 Just Save Your work has been saved! (Back to Admin Page) Submit Problem #3 for Grading Your Answer:1.20 0/2x 0.83 0.79 0/2x Your Mark: 0/2x
Answer #1

Present Value (PV) = 1000

Future Value (FV) = 2850

Time Period n = 11*12 = 132 months

Let effective monthly interest rate be r

=> FV = PV(1+r)n

=> 2850 = 1000 (1+r)132

=> r = 0.0080

Hence effective monthly rate = 0.80%